# Iterative Utility Judgment Framework via LLMs Inspired by Relevance in Philosophy

### v1

@@ 8 Conclusion
In this paper, we propose an Iterative utiliTy judgmEnt fraMework (ITEM) to enhance the utility judgment and QA performance of LLMs by interactions between the steps, inspired by Schutz’s philosophical discussion of relevance. Our framework achieves state-of-the-art performance in zero-shot scenarios, outperforming previous methods in utility judgments, ranking of topical relevance, and answer generation tasks. Our experiments also highlight the significance of dynamic interaction in achieving high performance and stability. Future directions include developing better fine-tuning strategies for utility judgments and creating end-to-end solutions for retrieval and utility judgments.
@@
In this paper, we proposed an iterative utility judgment framework inspired by relevance in philosophy. There are three primary limitations that should be acknowledged: (i) Our methods are applied in zero-shot scenarios without any training. The zero-shot method has inherent disadvantages, e.g., being sensitive to prompts and having unstable performance. The zero-shot approach itself does not enhance the LLMs’s inherent capability in utility judgments but rather employs strategies to improve performance on utility judgment tasks. Future research should explore designing more effective training methods, e.g., utilizing our iterative framework with self-evolution techniques Singh et al. (2023), to genuinely enhance the LLMs’s ability in utility judgments through training. (ii) The number of candidate passages in the search scenario is much larger than 20. The number of search results we assumed is too small. We need to continue to study utility judgments in large-scale scenarios in the future. (iii) Although our iterative framework is very effective, the number of iterations will increase the cost of calling large models. How to reduce the iteration cost is also something that needs further improvement in the future.
Ethics Statement
@@
In conducting this research, we have prioritized ethical considerations at every stage to ensure the responsible development and application of AI technologies. Our research does not rely on personally identifiable information or require manually annotated datasets. We firmly believe in the principles of open research and the scientific value of reproducibility. To this end, we have made all models, data, and code associated with our paper publicly available on GitHub. This transparency not only facilitates the verification of our findings by the community but also encourages the application of our methods in other contexts.
Appendix A Instruction Details
A.1 Instruction of Listwise and Pointwise Approaches
For the prompts of the NQ dataset using ChatGPT, we follow the setting of Zhang et al. (2024), otherwise, we use the following prompts. Following Sun et al. (2023), we input N passages using the form of multiple rounds of dialogue in the listwise approach. Considering the quality of the pseudo answer, we design a note, i.e., “The reference answer may not be the correct answer, but it provides a pattern of the correct answer.”, adding before judging. We do experiments on listwise and pointwise respectively, and find that listwise is not sensitive to this note, note has a great influence on pointwise, as shown in Table 6 Finally, we choose prompt as shown in Figure 9 and Figure 10.
| listwise | pointwise | |||
| Note | Note | Note | Note | |
| =1 | 53.56 | 53.91 | 49.44 | 48.34 |
| =2 | 55.30 | 55.65 | 51.09 | 49.44 |
| =3 | 54.57 | 54.86 | 51.74 | 50.02 |
| =4 | 53.87 | 55.02 | 51.90 | 50.30 |
| =5 | 53.99 | 54.86 | 51.65 | 50.60 |
Moreover, for the implicit answer, we consider “what information is necessary to answer the question” and “which information is necessary to answer the question” in the prompt. Experiments are shown in Table 7. Finally, we choose prompt as shown in Figure 15. For the experiments of ChatGPT API, we have the cost of approximately $248.
| TREC | WebAP | |||
| which | what | which | what | |
| =1 | 49.39 | 48.34 | 26.06 | 25.81 |
| =2 | 50.72 | 50.13 | 28.67 | 25.38 |
| =3 | 52.05 | 50.94 | 28.36 | 25.04 |
| =4 | 51.71 | 51.54 | 25.46 | 26.87 |
| =5 | 51.54 | 50.66 | 27.35 | 26.69 |
A.2 Instruction of the Ranking Approach
A.3 Instruction of Answer Generation
Li et al. (2023) utilize LLM to generate the missing information in the provided documents for the current question and then re-retrieve it as relevant feedback. Therefore, we have also designed two kinds of pseudo answers for utility judgments, i.e., (i) the explicit answer, which produces an answer based on the given information, and (ii) the implicit answer, which does not answer the question directly but gives the information necessary to answer the question. For “words/sentences”, the experiments found that the length of the explicit answers generated by Mistral using “sentences” is too long for factual questions, whereas Llama 3 and chatGPT use “sentences“ to generate answers of moderate length. So we only use “words“ on the TREC dataset and the NQ dataset using Mistral. The two instructions are shown in Figure 14 and Figure 15.
Appendix B Experiment Details
We utilized AI models such as ChatGPT to perform grammar corrections on portions of our paper.
B.1 Effect of Iteration Numbers
@@
The performance of different LLMs in different datasets is shown in Table 8, Table 9, Table 10, Table 11 and Table 12.
@@
| NDCG@5 | NDCG@10 | NDCG@20 | Utility-judgments F1 | |
| 1 | 71.29 / 70.57 | 72.90 / 72.69 | 84.56 / 84.08 | 52.34 / 42.02 |
| 2 | 72.27 / 72.86 | 75.16 / 75.48 | 85.76 / 86.09 | 53.40 / 42.10 |
| 3 | 73.24 / 74.27 | 75.53 / 75.78 | 86.59 / 86.80 | 56.27 / 44.18 |
| 4 | 73.58 / 75.35 | 75.52 / 76.83 | 86.46 / 87.23 | 56.07 / 44.68 |
| 5 | 73.12 / 74.61 | 74.75 / 76.20 | 85.95 / 86.82 | 57.82 / 44.25 |
B.2 Iteration Stop Conditions
We consider two stop conditions for ITEM, i.e., utility judgments and answer. The different performances of Mistral with different stop conditions are in Table 13.
| Iteration stop condiction | TREC | WebAP | ||
| Listwise | Pointwise | Listwise | Pointwise | |
| Utility (=1) | 53.78 | 49.44 | 29.76 | 29.10 |
| Utility (=2) | 55.65 | 51.09 | 29.83 | 29.88 |
| Utility (=3) | 54.86 | 51.74 | 31.65 | 30.50 |
| Utility (=4) | 55.02 | 51.90 | 28.96 | 31.00 |
| Utility (=5) | 54.86 | 51.65 | 31.86 | 30.85 |
| Answer 0.9 (=1) | 53.78 | 49.44 | 29.76 | 29.10 |
| Answer 0.9 (=2) | 55.65 | 50.45 | 29.83 | 29.88 |
| Answer 0.9 (=3) | 54.90 | 50.88 | 31.65 | 30.50 |
| Answer 0.9 (=4) | 55.02 | 51.23 | 28.96 | 31.00 |
| Answer 0.9 (=5) | 54.90 | 50.79 | 31.86 | 30.85 |
| Answer 0.8 (=1) | 53.78 | 49.44 | 29.76 | 29.10 |
| Answer 0.8 (=2) | 55.16 | 50.46 | 29.83 | 30.20 |
| Answer 0.8 (=3) | 54.35 | 50.85 | 31.65 | 30.81 |
| Answer 0.8 (=4) | 54.53 | 51.20 | 28.96 | 31.30 |
| Answer 0.8 (=5) | 54.35 | 50.76 | 31.86 | 31.35 |
| Answer 0.7 (=1) | 53.78 | 49.44 | 29.76 | 29.10 |
| Answer 0.7 (=2) | 55.37 | 50.00 | 29.62 | 29.95 |
| Answer 0.7 (=3) | 54.93 | 50.31 | 31.58 | 30.75 |
| Answer 0.7 (=4) | 55.11 | 50.51 | 29.20 | 31.11 |
| Answer 0.7 (=5) | 54.93 | 50.28 | 31.77 | 31.31 |
| Answer 0.6 (=1) | 53.78 | 49.44 | 29.76 | 29.10 |
| Answer 0.6 (=2) | 54.98 | 50.02 | 29.76 | 30.02 |
| Answer 0.6 (=3) | 55.34 | 50.27 | 32.28 | 30.55 |
| Answer 0.6 (=4) | 54.78 | 50.53 | 30.29 | 30.92 |
| Answer 0.6 (=5) | 55.34 | 50.23 | 32.14 | 31.05 |
| Answer 0.5 (=1) | 53.78 | 49.44 | 29.76 | 29.10 |
| Answer 0.5 (=2) | 55.30 | 49.98 | 31.42 | 30.15 |
| Answer 0.5 (=3) | 55.39 | 50.23 | 32.97 | 30.17 |
| Answer 0.5 (=4) | 55.10 | 50.42 | 31.73 | 29.92 |
| Answer 0.5 (=5) | 55.39 | 50.16 | 32.50 | 29.92 |
B.3 Quality of Utility Judgments
The retrieved data we use are a four-point scale, and we consider the highest level as having utility. To see the different points in our results for utility judgments, we analyzed the various ranks as a percentage of the results on TREC for mistral in Table 14.
@@ B.3 Quality of Utility Judgments
From Table 14, we can see that almost 70% of the results of utility judgments are very relevant to the question.
B.4 values in ITEM-Ar
@@
Appendix C Case Study
Appendix D Annotations in NQ Dataset
To evaluate the performance of utility judgments under different values of , we relabeled the retrieved passages following Stelmakh et al. (2022). Specifically, we used a QA model based on RoBERTa Liu et al. (2019) to extract answers from the retrieved passages. A passage is considered utility if the extracted answers match the ground-truth answers. The advantage of this labeling method is its relatively high accuracy, while its disadvantage is the potentially low recall rate.
Appendix E Datasets
Detailed statistics of the experimental datasets are shown in Table 16.
@@ Appendix E Datasets
Appendix F Answer Passage Retrieval

### v2

@@ 9 Conclusion
In this paper, we propose an Iterative utiliTy judgmEnt fraMework (ITEM) to enhance the utility judgment and QA performance of LLMs by interactions between the steps, inspired by Schutz’s philosophical discussion of relevance. This is a unified framework of iterative RAG with an emphasis on utility. Our framework achieves state-of-the-art performance in zero-shot scenarios, outperforming previous methods in utility judgments, ranking of topical relevance, and answer generation tasks, indicating that the cognitive process of LLMs on a specific topic can also be improved by a similar process. Our experiments also highlight the significance of dynamic interaction in achieving high performance and stability. Future directions include developing better fine-tuning strategies for utility judgments and creating end-to-end solutions for RAG.
@@
There are two primary limitations that should be acknowledged: (i) Our methods are applied in zero-shot scenarios without any training. The zero-shot approach itself does not enhance the LLMs’s inherent capability in utility judgments but rather employs strategies to improve performance on utility judgment tasks. Future research should explore designing more effective training methods, e.g., utilizing our iterative framework with self-evolution techniques Singh et al. (2023), to genuinely enhance the LLMs’s ability in utility judgments through training. (ii) The number of candidate passages in the search scenario is much larger than 20. The number of search results we assumed is too small. We need to continue to study utility judgments in large-scale scenarios in the future.
10 Ethics Statement
@@
Our research does not rely on personally identifiable information. All datasets and models used in our paper are publicly available and have been widely adopted by researchers. We firmly believe in the principles of open research and the scientific value of reproducibility. To this end, we have made all data, and code associated with our paper publicly available on GitHub. This transparency not only facilitates the verification of our findings by the community but also encourages the application of our methods in other contexts.
Appendix A Experiment Details
A.1 Effect of Iteration Numbers
@@
The precision, recall, and F1 performance of different LLMs on different datasets with different iteration numbers is shown in Table 6, Table 7, Table 8, and Table 9.
@@
A.2 Quality of Utility Judgments
The relevance labels of TREC DL are of a four-point scale, and we consider the highest level as having utility. To see the utility judgment performance when we consider lower grades to have utility, we measure the precision of utility judgments of Mistral on TREC DL when passages of different grades are treated as positive in Table 10.
@@ A.2 Quality of Utility Judgments
We can see that almost 70% of the results of positive utility judgments are highly relevant to the question.
A.3 values in ITEM-Ar
@@
A.4 ITEM-ARr
We evaluate two ranking performances of ITEM-ARr during the same loop, with the experimental results shown in Table 12. We find that under the ITEM-ARr framework, relevance ranking and utility ranking are both improved, and utility ranking performance is generally better than relevance ranking. However, as seen in Table 3 and Table 12, performing ranking twice in the same iteration may not yield better ranking results than performing utility ranking once in the iteration.
| NDCG@5 | NDCG@10 | NDCG@20 | Utility-F1 | |
| 1 | 71.29 / 72.77 | 72.90 / 74.96 | 84.56 / 85.75 | 43.13 |
| 2 | 72.54 / 70.99 | 74.81 / 73.76 | 85.77 / 85.28 | 40.21 |
| 3 | 72.07 / 74.14 | 74.14 / 76.63 | 85.53 / 86.57 | 45.67 |
| 4 | 71.02 / 71.06 | 74.30 / 74.03 | 85.09 / 85.16 | 43.82 |
| 5 | 72.26 / 70.12 | 75.83 / 72.59 | 85.88 / 84.77 | 44.10 |
Appendix B Case Study
We show two cases on the TREC dataset in Table 9.
Appendix C Datasets and Evaluation
@@
| GTI-NQ | 10 | 100 | 1863 | 1.0 |
Detailed statistics of the experimental datasets are shown in Table 13. We use the official tool for evaluation of NDCG.
Appendix D Answer Passage Retrieval
@@
Appendix F Retrievers
We use two representative retrievers to gather candidate passages in for utility judgments. Following with previous works Zhang et al. (2024); Sun et al. (2023), we employ RocketQAv2 Ren et al. (2021) and BM25 Robertson et al. (2009) for the NQ dataset and retrieval datasets(i.e., TREC DL and WebAP datasets), respectively. Based on the retrieval results to build the we have two settings: (i) For TREC DL and WebAP datasets, we select the top-20 BM25 retrieval results. If these do not include passages with utility, we replaced the last one with a utility-annotated passage. (ii) For the NQ dataset, we use the top-10 dense retrieval results to form the candidate list , following the GTU setting of Zhang et al. (2024).
Appendix G Inference Efficiency
Table 14 shows more analysis of Mistral on the TREC DL dataset using the listwise input form. The temperature of the LLMs is set to 0 during the generation process, and we used a single run.
| Methods | #IT/Q | #OT/Q | #RT(s)/Q |
| Vanilla | 2507 | 23 | 0.49 |
| UJ-ExpA | 2529 | 59 | 0.94 |
| UJ-ImpA | 2533 | 41 | 0.71 |
| 5-sampling | 12647 | 296 | 4.72 |
| ITEM-As(m=1) | 4628 | 47 | 0.96 |
| ITEM-As(m=3) | 10603 | 154 | 2.13 |
| ITEM-ARs(m=1) | 7107 | 211 | 3.15 |
| ITEM-ARs(m=3) | 18224 | 624 | 8.61 |
Due to the iterations, the average input token length of our methods is relatively large. The cost of ITEM-As is about 0.5 times that of the 5-sampling, and ITEM-ARs is about 1.5 times the cost of 5-sampling. Our framework provides a way of automatically obtaining high-quality labeled data for each task in RAG. These annotations can be used to train regular RAG models. The cost should be worth it (much better than -sampling).
Appendix H Instruction Details
H.1 Instruction of Listwise and Pointwise Approaches
For the prompts of the NQ dataset using ChatGPT, we follow the setting of Zhang et al. (2024), otherwise, we use the following prompts. Following Sun et al. (2023), we input N passages using the form of multiple rounds of dialogue in the listwise approach. The prompts we used in our experiments are shown in Figure 10 and Figure 11.
H.2 Instruction of the Ranking Approach
H.3 Instruction of Answer Generation
Li et al. (2023) utilize LLM to generate the missing information in the provided documents for the current question and then re-retrieve it as relevant feedback. Therefore, we have also designed two kinds of pseudo answers for utility judgments, i.e., (i) the explicit answer, which produces an answer based on the given information, and (ii) the implicit answer, which does not answer the question directly but gives the information necessary to answer the question. The two instructions are shown in Figure 15 and Figure 16.